# The Curse of Depth in Large Language Models

### v2

@@ Abstract
In this paper, we introduce the Curse of Depth, a concept that highlights, explains, and addresses the recent observation in modern Large Language Models (LLMs) where nearly half of the layers are less effective than expected. We first confirm the wide existence of this phenomenon across the most popular families of LLMs such as Llama, Mistral, DeepSeek, and Qwen. Our analysis, theoretically and empirically, identifies that the underlying reason for the ineffectiveness of deep layers in LLMs is the widespread usage of Pre-Layer Normalization (Pre-LN). While Pre-LN stabilizes the training of Transformer LLMs, its output variance exponentially grows with the model depth, which undesirably causes the derivative of the deep Transformer blocks to be an identity matrix, and therefore barely contributes to the training. To resolve this training pitfall, we propose LayerNorm Scaling (LNS), which scales the variance of output of the layer normalization inversely by the square root of its depth. This simple modification mitigates the output variance explosion of deeper Transformer layers, improving their contribution. Across a wide range of model sizes (130M to 7B), our experiments show that LNS consistently outperforms previous normalization and scaling techniques in enhancing LLM pre-training performance. Moreover, this improvement seamlessly carries over to supervised fine-tuning. All these gains can be attributed to the fact that LayerNorm Scaling enables deeper layers to contribute more effectively during training. Our code is available at LayerNorm-Scaling.
@@
Figure 1: Left: Schematic diagrams of (a) Pre-LN and (b) LayerNorm Scaling. LayerNorm Scaling applies a scaling factor inversely proportional to the square root of the layer index $\ell$, preventing excessive variance growth. Right: Language modeling loss of scaling up parameter count up to 7B. All models are trained for 20B tokens using OLMo (Groeneveld et al., 2024).
@@ 1 Introduction
Recent studies reveal that the deeper layers (Transformer blocks) in modern LLMs tend to be less effective than the earlier ones (Yin et al., 2024; Gromov et al., 2024; Men et al., 2024; Li et al., 2024b). On the one hand, this interesting observation provides an effective indicator for LLM compression. For instance, we can compress deeper layers significantly more (Yin et al., 2024; Lu et al., 2024; Dumitru et al., 2024) to achieve high compression ratios. Even more aggressively, entire deep layers can be pruned completely without compromising performance (Muralidharan et al., 2024; Siddiqui et al., 2024).
On the other hand, having many layers ineffective is undesirable as modern LLMs are extremely resource-intensive to train, often requiring thousands of GPUs trained for multiple months, let alone the labor used for data curation and administration Achiam et al. (2023); Touvron et al. (2023). Ideally, we want all layers in a model to be well-trained, with sufficient diversity in features from layer to layer, to maximize the utility of resources (Li et al., 2024b). The existence of ill-trained layers suggests that there must be something off with current LLM paradigms. Addressing such limitations is a pressing need for the community to avoid the waste of valuable resources, as new versions of LLMs are usually trained with their previous computing paradigm which results in ineffective layers.
@@
Empirical Evidence of CoD. To demonstrate that CoD is a common phenomenon across prominent LLM families, we perform layer pruning experiments on Qwen3, LLaMA2, and DeepSeek. Specifically, we prune one layer at a time, without any fine-tuning, and directly evaluate the resulting pruned models on the MMLU benchmark (Hendrycks et al., 2021), as shown in Figure 2. Key findings: (1) Most models, including the latest Qwen3, exhibit surprising resilience to the removal of deeper layers; (2) The number of layers that can be removed without causing significant performance drop increases with model size; (3) Representations in deeper layers are significantly more similar to each other than those in earlier layers.
Identifying the Root Cause of CoD. We theoretically and empirically identify the root cause of CoD as the use of Pre-Layer Normalization (Pre-LN) (Baevski and Auli, 2019; Dai et al., 2019), which normalizes layer inputs before applying the main computations, such as attention or feedforward operations, rather than after. Specifically, while stabilizing training, we observe that the output variance of Pre-LN accumulates significantly with layer depth as shown in Figure 4, causing the derivatives of deep Pre-LN layers to approach an identity matrix. This behavior prevents these layers from introducing meaningful transformations, leading to diminished representation learning.
Mitigating CoD through LayerNorm Scaling. We propose LayerNorm Scaling (LNS), which scales the output of Layer Normalization by the square root of the depth $\frac{1}{\sqrt{l}}$. LayerNorm Scaling effectively scales down the output variance across layers of Pre-LN. LNS consistently delivers better pre-training performance than existing normalization and scaling techniques across various model sizes from 130M to 7B. Unlike previous LayerNorm variants (Li et al., 2024b; Liu et al., 2020), LayerNorm Scaling is simple to implement, requires no hyperparameter tuning, and introduces no additional parameters during training. [^1] We found that combining LNS with Scaled Initialization (Groeneveld et al., 2024; Radford et al., 2019; Shoeybi et al., 2020) diminishes the effectiveness of LNS. Therefore, we recommend removing the latter when applying LNS. Furthermore, we show that the model pre-trained with LayerNorm Scaling achieves better performance on downstream tasks in self-supervised fine-tuning, all thanks to the more diverse feature representations learned in deep layers.
@@ 2 Empirical Evidence of the Curse of Depth
To empirically analyze the impact of layer normalization on the Curse of Depth in LLMs, we conduct a series of evaluations inspired by Li et al. (2024b), to compare Pre-LN and Post-LN models.
@@ 2.1 Open-weight Large-scale LLMs
Models: To verify this, we empirically quantify the contribution of individual layers to overall model performance across a diverse set of LLMs, including Qwen3 (Team, 2025), LLaMA2 (Touvron et al., 2023), DeepSeek (Bi et al., 2024), and BERT-Large (Devlin, 2019). These models were chosen to ensure architectural and application diversity. BERT-Large represents a Post-LN model, whereas the rest are Pre-LN-based. This selection enables a comprehensive evaluation of the effects of layer normalization across varying architectures and model scales.
Evaluation Metric: To empirically assess the impact of deeper layers in LLMs, we adopt two metrics, Performance Drop and Angular Distance, inspired by Gromov et al. (2024); Li et al. (2024b).
Performance Drop $\Delta P^{(\ell)}$ quantifies the importance of each layer by measuring the performance change after its removal. A smaller $\Delta P^{(\ell)}$ indicates that the pruned layer contributes less to the model’s overall performance. For BERT-Large, we evaluate using the SQuAD v1.1 dataset (Rajpurkar, 2016), while for other models, we use MMLU (Hendrycks et al., 2021), a standard benchmark for multi-task language understanding.
@@
where $\|\cdot\|_{2}$ denotes the $L^{2}$-norm. To reduce variance, we report the average distance over 256K tokens sampled from the C4 dataset. Smaller values of $d(x^{\ell},x^{\ell+n})$ indicate higher similarity between the two representations, suggesting limited transformation. Such layers can be considered redundant, as their removal minimally impacts the model’s internal representations. Ideally, each layer should introduce meaningful representational shifts to fully leverage the model’s capacity (Yang et al., 2023; Gromov et al., 2024).
@@
Figure 4: Layerwise output variance. This figure compares the output variance across various layers for different setups: (1) Pre-LN; (2) Pre-LN with Scaled Initialization (Shoeybi et al., 2020; Radford et al., 2019); and (3) LayerNorm Scaling. The experiments are conducted on the LLaM-130M model trained for 10,000 steps. The proposed LayerNorm Scaling effectively controls the variance across layers.
@@ 2.2 In-house Small-scale LLaMa-130M
To eliminate the influence of other confounding variables, we train two LLaMA-130M models from scratch that differ only in their Layer Normalization, thereby clearly distinguishing Post-LN from Pre-LN, following Li et al. (2024b). The results are illustrated in Figure 3.
@@ 3 Analysis of the Curse of Depth
Preliminaries. This paper primarily focuses on Pre-LN Transformer (Baevski and Auli, 2019; Dai et al., 2019). Let $x_{\ell}\in\mathbb{R}^{d}$ be the input vector at the $\ell$-th layer of Transformer, where $d$ denotes the feature dimension of each layer. For simplicity, we assume all layers to have the same dimension $d$. The layer output $y$ is calculated as follows:
@@ 3.1 Pre-LN Transformers
Let $x_{\ell}$ and $x^{\prime}_{\ell}$ denote the input and intermediate vectors of the $\ell$-th layer. Moreover, let $W_{\ell}$ denote the model parameter matrix at the $\ell$-th layer. We assume that, for all layers, $x_{\ell}$, $x^{\prime}_{\ell}$, and $W_{\ell}$ follow normal and independent distributions with mean $\mu=0$.
@@
Based on Assumption 3.1 and the work of Takase et al. (2023b), we obtain the following:
@@
The detailed description of $A$ and $B$, as well as the complete proof, are provided in Appendix A.2. From Theorem 3.3, we observe that when the variance grows exponentially, as the number of layers $L\to\infty$, the norm $\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}$ is bounded above by a fixed constant $M$. This result implies that even an infinitely deep Transformer remains stable, and by the Weierstrass Theorem, the network is guaranteed to converge. Consequently, this implies that for very large $L$, deeper layers behave nearly as an identity map from $x_{\ell}$ to $y_{\ell}$, thereby limiting the model’s expressivity and hindering its ability to learn meaningful transformations. This outcome is undesirable, therefore, we would instead prefer the variance to increase more gradually—e.g., linearly—so that $\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}$ exhibits linear growth. This observation highlights the necessity of appropriate variance control mechanisms, such as scaling strategies, to prevent excessive identity mappings and enhance network depth utilization.
@@ 4 LayerNorm Scaling (LNS)
where $\ell\in\{1,2,\dots,L\}$. This scaling prevents excessive variance growth with depth, addressing a key limitation of Pre-LN. Unlike Mix-LN, which stabilizes gradients in deeper layers but suffers from training instability caused by Post-LN (Nguyen and Salazar, 2019; Wang et al., 2024), LayerNorm Scaling preserves the stability advantages of Pre-LN while enhancing the contribution of deeper layers to representation learning. Applying LayerNorm Scaling leads to a notable reduction of layerwise output variance as shown in Figure 4, resulting in a lower training loss. Moreover, compared with previous LayerNorm variants (Li et al., 2024b; Liu et al., 2020), LayerNorm Scaling is hyperparameter-free, easy to implement, and does not introduce additional learnable parameters, making it computationally efficient and readily applicable to existing Transformer architectures.
@@ 4.1 Theoretical Analysis of LayerNorm Scaling
where $\epsilon$ is a small number with $11/2\leq\epsilon<$ 1.
@@
In addition to making the deeper layers more effective, our variance-scaling approach can also reduce sudden spikes in the loss landscape during training. Based on Takase et al. (2023b)’s work, We formalize this in the following theorem Theorem 4.3, which gives a rigorous upper bound on the gradient norm with respect to the attention parameters.
@@ 5.1 LLM Pre-training
To evaluate the effectiveness of LayerNorm Scaling, we follow the experimental setup of Li et al. (2024b), using the identical model configurations and training conditions to compare LNS with widely used normalization techniques, including Post-LN (Nguyen and Salazar, 2019), DeepNorm (Wang et al., 2024), and Pre-LN (Dai et al., 2019). In line with Lialin et al. (2023) and Zhao et al. (2024), we conduct experiments using LLaMA-based architectures with model sizes of 130M, 250M, 350M, and 1B parameters.
@@
Post-LN (Ba, 2016) 26.95 1409.79 1368.33 1390.75
DeepNorm (Wang et al., 2024) 27.17 22.77 1362.59 1409.08
Mix-LN (Li et al., 2024b) 26.07 21.39 1363.21 1414.78
Pre-LN (Baevski and Auli, 2019) 26.73 21.92 19.58 17.02
@@
The architecture incorporates RMSNorm (Shazeer, 2020) and SwiGLU activations (Zhang and Sennrich, 2019), which are applied consistently across all model sizes and normalization methods. For optimization, we use the Adam optimizer (Kingma, 2015) and adopt size-specific learning rates: $1\times 10^{-3}$ for models up to 350M parameters, and $5\times 10^{-4}$ for the 1B parameter model. All models share the same architecture, hyperparameters, and training schedule, with the only difference being the choice of normalization method. Unlike Mix-LN (Li et al., 2024b), which introduces an additional hyperparameter $\alpha$ manually set to 0.25, LayerNorm Scaling requires no extra hyperparameters, making it simpler to implement. Table 1 shows that LayerNorm Scaling consistently outperforms other normalization methods across different model sizes. While DeepNorm performs comparably to Pre-LN on smaller models, it struggles with larger architectures like LLaMA-1B, showing signs of instability and divergence in loss values. Similarly, Mix-LN outperforms Pre-LN in smaller models but faces convergence issues with LLaMA-350M, indicating its sensitivity to architecture design and hyperparameter tuning due to the introduction of Post-LN. Notably, Mix-LN was originally evaluated on LLaMA-1B with 50K steps (Li et al., 2024b), while our setting extends training to 100K steps, where Mix-LN fails to converge, highlighting its instability in large-scale settings caused by the usage of Post-LN.

### v3

@@ Abstract
In this paper, we introduce the Curse of Depth, a concept that highlights, explains, and addresses the recent observation in modern Large Language Models (LLMs) where nearly half of the layers are less effective than expected. We first confirm the wide existence of this phenomenon across the most popular families of LLMs such as Llama, Mistral, DeepSeek, and Qwen. Our analysis, theoretically and empirically, identifies that the underlying reason for the ineffectiveness of deep layers in LLMs is the widespread usage of Pre-Layer Normalization (Pre-LN). While Pre-LN stabilizes the training of Transformer LLMs, its output variance exponentially grows with the model depth, which undesirably causes the derivative of the deep Transformer blocks to be an identity matrix, and therefore barely contributes to the training. To resolve this training pitfall, we propose LayerNorm Scaling (LNS), which scales the variance of output of the layer normalization inversely by the square root of its depth. [^1] We found that combining LNS with Scaled Initialization (groeneveld2024olmo; radford2019language; shoeybi2019megatron) diminishes the effectiveness of LNS. Therefore, we recommend removing the latter when applying LNS. This simple modification mitigates the output variance explosion of deeper Transformer layers, improving their contribution. Across a wide range of model sizes (130M to 7B), our experiments show that LNS consistently outperforms previous normalization and scaling techniques in enhancing LLM pre-training performance. Moreover, this improvement seamlessly carries over to supervised fine-tuning. All these gains can be attributed to the fact that LayerNorm Scaling enables deeper layers to contribute more effectively during training. Our code is available at LayerNorm-Scaling.
@@
Figure 1: Left: Schematic diagrams of (a) Pre-LN and (b) LayerNorm Scaling. LayerNorm Scaling applies a scaling factor inversely proportional to the square root of the layer index $\ell$, preventing excessive variance growth. Right: Language modeling loss of scaling up parameter count up to 7B. All models are trained for 20B tokens using OLMo (groeneveld2024olmo).
@@ 1 Introduction
Recent studies reveal that the deeper layers (Transformer blocks) in modern LLMs tend to be less effective than the earlier ones (yin2023outlier; gromov2024unreasonable; men2024shortgpt; li2024mix). On the one hand, this interesting observation provides an effective indicator for LLM compression. For instance, we can compress deeper layers significantly more (yin2023outlier; lu2024alphapruning; dumitru2024layer) to achieve high compression ratios. Even more aggressively, entire deep layers can be pruned completely without compromising performance (muralidharan2024compact; siddiqui2024deeper).
On the other hand, having many layers ineffective is undesirable as modern LLMs are extremely resource-intensive to train, often requiring thousands of GPUs trained for multiple months, let alone the labor used for data curation and administration achiam2023gpt; touvron2023llama. Ideally, we want all layers in a model to be well-trained, with sufficient diversity in features from layer to layer, to maximize the utility of resources (li2024mix). The existence of ill-trained layers suggests that there must be something off with current LLM paradigms. Addressing such limitations is a pressing need for the community to avoid the waste of valuable resources, as new versions of LLMs are usually trained with their previous computing paradigm which results in ineffective layers.
@@
Empirical Evidence of CoD. To demonstrate that CoD is a common phenomenon across prominent LLM families, we perform layer pruning experiments on Qwen3, LLaMA2, and DeepSeek. Specifically, we prune one layer at a time, without any fine-tuning, and directly evaluate the resulting pruned models on the MMLU benchmark (hendrycks2020measuring), as shown in Figure 2. Key findings: (1) Most models, including the latest Qwen3, exhibit surprising resilience to the removal of deeper layers; (2) The number of layers that can be removed without causing significant performance drop increases with model size; (3) Representations in deeper layers are significantly more similar to each other than those in earlier layers.
Identifying the Root Cause of CoD. We theoretically and empirically identify the root cause of CoD as the use of Pre-Layer Normalization (Pre-LN) (baevski2018adaptive; dai2019transformer), which normalizes layer inputs before applying the main computations, such as attention or feedforward operations, rather than after. Specifically, while stabilizing training, we observe that the output variance of Pre-LN accumulates significantly with layer depth as shown in Figure 4, causing the derivatives of deep Pre-LN layers to approach an identity matrix. This behavior prevents these layers from introducing meaningful transformations, leading to diminished representation learning.
Mitigating CoD through LayerNorm Scaling. We propose LayerNorm Scaling (LNS), which scales the output of Layer Normalization by the square root of the depth $\frac{1}{\sqrt{l}}$. LayerNorm Scaling effectively scales down the output variance across layers of Pre-LN. LNS consistently delivers better pre-training performance than existing normalization and scaling techniques across various model sizes from 130M to 7B. Unlike previous LayerNorm variants (li2024mix; liu2020understanding), LayerNorm Scaling is simple to implement, requires no hyperparameter tuning, and introduces no additional parameters during training. Furthermore, we show that the model pre-trained with LayerNorm Scaling achieves better performance on downstream tasks in self-supervised fine-tuning, all thanks to the more diverse feature representations learned in deep layers.
@@ 2 Empirical Evidence of the Curse of Depth
To empirically analyze the impact of layer normalization on the Curse of Depth in LLMs, we conduct a series of evaluations inspired by li2024mix, to compare Pre-LN and Post-LN models.
@@ 2.1 Open-weight Large-scale LLMs
Models: To verify this, we empirically quantify the contribution of individual layers to overall model performance across a diverse set of LLMs, including Qwen3 (qwen3_blog), LLaMA2 (touvron2023llama), DeepSeek (bi2024deepseek), and BERT-Large (devlin2018bert). These models were chosen to ensure architectural and application diversity. BERT-Large represents a Post-LN model, whereas the rest are Pre-LN-based. This selection enables a comprehensive evaluation of the effects of layer normalization across varying architectures and model scales.
Evaluation Metric: To empirically assess the impact of deeper layers in LLMs, we adopt two metrics, Performance Drop and Angular Distance, inspired by gromov2024unreasonable; li2024mix.
Performance Drop $\Delta P^{(\ell)}$ quantifies the importance of each layer by measuring the performance change after its removal. A smaller $\Delta P^{(\ell)}$ indicates that the pruned layer contributes less to the model’s overall performance. For BERT-Large, we evaluate using the SQuAD v1.1 dataset (rajpurkar2016squad), while for other models, we use MMLU (hendrycks2020measuring), a standard benchmark for multi-task language understanding.
@@
where $\|\cdot\|_{2}$ denotes the $L^{2}$-norm. To reduce variance, we report the average distance over 256K tokens sampled from the C4 dataset. Smaller values of $d(x^{\ell},x^{\ell+n})$ indicate higher similarity between the two representations, suggesting limited transformation. Such layers can be considered redundant, as their removal minimally impacts the model’s internal representations. Ideally, each layer should introduce meaningful representational shifts to fully leverage the model’s capacity (yang2023tensor; gromov2024unreasonable).
@@
Figure 4: Layerwise output variance. This figure compares the output variance across various layers for different setups: (1) Pre-LN; (2) Pre-LN with Scaled Initialization (shoeybi2019megatron; radford2019language); and (3) LayerNorm Scaling. The experiments are conducted on the LLaM-130M model trained for 10,000 steps. The proposed LayerNorm Scaling effectively controls the variance across layers.
@@ 2.2 In-house Small-scale LLaMa-130M
To eliminate the influence of other confounding variables, we train two LLaMA-130M models from scratch that differ only in their Layer Normalization, thereby clearly distinguishing Post-LN from Pre-LN, following li2024mix. The results are illustrated in Figure 3.
@@ 3 Analysis of the Curse of Depth
Preliminaries. This paper primarily focuses on Pre-LN Transformer (baevski2018adaptive; dai2019transformer). Let $x_{\ell}\in\mathbb{R}^{d}$ be the input vector at the $\ell$-th layer of Transformer, where $d$ denotes the feature dimension of each layer. For simplicity, we assume all layers to have the same dimension $d$. The layer output $y$ is calculated as follows:
@@ 3.1 Pre-LN Transformers
Let $x_{\ell}$ and $x^{\prime}_{\ell}$ denote the input and intermediate vectors of the $\ell$-th layer. Moreover, let $W_{\ell}$ denote the model parameter matrix at the $\ell$-th layer. We assume that, for all layers, $_{\ell}, xx^{\prime}_{\ell}$, and $W_{\ell}$ follow normal and independent distributions with mean $\mu=0$.
@@
Based on Assumption 3.1 and the work of takase2023spike, we obtain the following:
@@
The detailed description of $A$ and $B$, as well as the complete proof, are provided in Appendix A.2. From Theorem 3.3, we observe that when the variance grows exponentially, as the number of layers $L\to\infty$, the norm $\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}$ is bounded above by a fixed constant $M$. This result implies that even an infinitely deep Transformer remains stable, and by the Weierstrass Theorem, the network is guaranteed to converge. Consequently, this implies that for very large $L$, deeper layers behave nearly as an identity map from $x_{\ell}$ to $y_{\ell}$, thereby limiting the model’s expressivity and hindering its ability to learn meaningful transformations. This phenomenon is empirically illustrated in Figure 5, where we visualize the LayerNorm Jacobians of a pre-trained LLaMA2-7B model across depth, revealing a clear collapse toward identity mappings in deeper layers. This outcome is undesirable, therefore, we would instead prefer the variance to increase more gradually—e.g., linearly—so that $\left\|\frac{\partial y_{L}}{\partial x_{1}}\right\|_{2}$ exhibits linear growth. This observation highlights the necessity of appropriate variance control mechanisms, such as scaling strategies, to prevent excessive identity mappings and enhance network depth utilization.
Figure 5: Visualization of the LayerNorm Jacobian matrices across different layers of a pre-trained LLaMA2-7B model. Each heatmap shows the token-averaged Jacobian at a specific layer. As depth increases, the Jacobians exhibit a pronounced diagonal dominance with vanishing off-diagonal entries, indicating that deep LayerNorm blocks increasingly approximate identity mappings.
@@ 4 LayerNorm Scaling (LNS)
where $\ell\in\{1,2,\dots,L\}$. This scaling prevents excessive variance growth with depth, addressing a key limitation of Pre-LN. Unlike Mix-LN, which stabilizes gradients in deeper layers but suffers from training instability caused by Post-LN (nguyen2019transformers; wang2024deepnet), LayerNorm Scaling preserves the stability advantages of Pre-LN while enhancing the contribution of deeper layers to representation learning. Applying LayerNorm Scaling leads to a notable reduction of layerwise output variance as shown in Figure 4, resulting in a lower training loss. Moreover, compared with previous LayerNorm variants (li2024mix; liu2020understanding), LayerNorm Scaling is hyperparameter-free, easy to implement, and does not introduce additional learnable parameters, making it computationally efficient and readily applicable to existing Transformer architectures.
@@ 4.1 Theoretical Analysis of LayerNorm Scaling
where $\epsilon$ is a small number with $1/2\leq\epsilon<1$.
@@
In addition to making the deeper layers more effective, our variance-scaling approach can also reduce sudden spikes in the loss landscape during training. Based on takase2023spike’s work, We formalize this in the following theorem Theorem 4.3, which gives a rigorous upper bound on the gradient norm with respect to the attention parameters.
@@ 5.1 LLM Pre-training
To evaluate the effectiveness of LayerNorm Scaling, we follow the experimental setup of li2024mix, using the identical model configurations and training conditions to compare LNS with widely used normalization techniques, including Post-LN (nguyen2019transformers), DeepNorm (wang2024deepnet), and Pre-LN (dai2019transformer). In line with lialin2023relora and zhao2024galore, we conduct experiments using LLaMA-based architectures with model sizes of 130M, 250M, 350M, and 1B parameters.
@@
Post-LN (ba2016layer) 26.95 1409.79 1368.33 1390.75
DeepNorm (wang2024deepnet) 27.17 22.77 1362.59 1409.08
Mix-LN (li2024mix) 26.07 21.39 1363.21 1414.78
Pre-LN (baevski2018adaptive) 26.73 21.92 19.58 17.02
@@
The architecture incorporates RMSNorm (shazeer2020glu) and SwiGLU activations (zhang2019root), which are applied consistently across all model sizes and normalization methods. For optimization, we use the Adam optimizer (kingma2014adam) and adopt size-specific learning rates: $1\times 10^{-3}$ for models up to 350M parameters, and $5\times 10^{-4}$ for the 1B parameter model. All models share the same architecture, hyperparameters, and training schedule, with the only difference being the choice of normalization method. Unlike Mix-LN (li2024mix), which introduces an additional hyperparameter $\alpha$ manually set to 0.25, LayerNorm Scaling requires no extra hyperparameters, making it simpler to implement. Table 1 shows that LayerNorm Scaling consistently outperforms other normalization methods across different model sizes. While DeepNorm performs comparably to Pre-LN on smaller models, it struggles with larger architectures like LLaMA-1B, showing signs of instability and divergence in loss values. Similarly, Mix-LN outperforms Pre-LN in smaller models but faces convergence issues with LLaMA-350M, indicating its sensitivity to architecture design and hyperparameter tuning due to the introduction of Post-LN. Notably, Mix-LN was originally evaluated on LLaMA-1B with 50K steps (li2024mix), while our setting extends training to 100K steps, where Mix-LN fails to converge, highlighting its instability in large-scale settings caused by the usage of Post-LN.